OpenStudy (lilcathyx):

Using the following equation, find the center and radius of the circle. You must show all work and calculations to receive credit. x^2 − 4x + y^2 + 8y = −4

OpenStudy (lilcathyx):

i have no clue on how to do this. & i don't know the complete square thingy.

OpenStudy (lilcathyx):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

focus on just x^2 - 4x

OpenStudy (lilcathyx):

ok

jimthompson5910 (jim_thompson5910):

is x^2 - 4x a perfect square?

OpenStudy (lilcathyx):

yes ?

jimthompson5910 (jim_thompson5910):

so you can factor x^2 - 4x into something of the form (a+b)^2 ?

OpenStudy (lilcathyx):

yes

jimthompson5910 (jim_thompson5910):

what does it factor to?

OpenStudy (lilcathyx):

(x - 2) ^ 2 ?

jimthompson5910 (jim_thompson5910):

let's check that

jimthompson5910 (jim_thompson5910):

(x - 2) ^ 2 = (x-2)(x-2) (x - 2) ^ 2 = x^2 - 2x - 2x + 4 (x - 2) ^ 2 = x^2 - 4x + 4

jimthompson5910 (jim_thompson5910):

so x^2 - 4x isn't a perfect square because it's missing that +4 at the end IF it had the +4 at the end, then x^2 - 4x + 4 is definitely a perfect square since it factors to (x-2)^2

jimthompson5910 (jim_thompson5910):

what we can do is add 4 to both sides to fix that issue x^2 - 4x + y^2 + 8y = -4 x^2 - 4x + y^2 + 8y+4 = -4+4 ... Add 4 to both sides. (x^2 - 4x + 4) + y^2 + 8y = 0

jimthompson5910 (jim_thompson5910):

we can then factor the stuff in the parenthesis to get (x-2)^2 so we go from (x^2 - 4x + 4) + y^2 + 8y = 0 to (x-2)^2 + y^2 + 8y = 0 making sense?

OpenStudy (lilcathyx):

a lil

OpenStudy (lilcathyx):

so now we do that to y^2 +8 = 0? @jim_thompson5910

OpenStudy (lilcathyx):

why not (x-2)^2 + 4 and then so on.

jimthompson5910 (jim_thompson5910):

we have (x-2)^2 + y^2 + 8y = 0

jimthompson5910 (jim_thompson5910):

y^2 + 8y is NOT a perfect square (y+4)^2 = y^2 + 8y + 16 so we have to add 16 to both sides to allow us to factor (x-2)^2 + y^2 + 8y = 0 (x-2)^2 + y^2 + 8y `+16` = 0 `+16` (x-2)^2 + (y^2 + 8y +16) = 16 (x-2)^2 + (y + 4)^2 = 16

jimthompson5910 (jim_thompson5910):

so in short, we go from x^2 - 4x + y^2 + 8y = -4 to (x-2)^2 + y^2 + 8y = 0 and then that turns into (x-2)^2 + (y + 4)^2 = 16 this is just a shorter version of the steps to give a summary in a way. Of course, look over each step carefully to make sure you know what's going on.

OpenStudy (lilcathyx):

okk. im going to look at every step you did.

jimthompson5910 (jim_thompson5910):

ok don't hesitate to let me know where you get stuck if you do at all